# Hyper compressed fine-tuning of large foundation models with Quantum Inspired Adapters

### v1

@@ 4.3 Adapter Configurations
To thoroughly evaluate the effectiveness of our Quantum-Inspired Adapters, we systematically varied several key aspects of their configuration, as discussed above. Our experimentation focused on different combinations of compound matrices based on Hamming-weights, the types of operations applied to these compounds, the enforcement of orthogonality, and the strategy for parameter sharing across adapter blocks.
@@
$\mathcal{C}:=\{\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{C}_{3}\},\ \ \mathcal{P}(\mathcal{C})^{\oplus 3}:=\{\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{C}_{3},\mathcal{C}_{1}\oplus\mathcal{C}_{2},\mathcal{C}_{2}\oplus\mathcal{C}_{3},\mathcal{C}_{2}\oplus\mathcal{C}_{3},\mathcal{C}_{1}\oplus\mathcal{C}_{2}\oplus\mathcal{C}_{3}\}$ (14)
Note that this notation is slightly obfuscating - for a fixed pre-trained matrix and block size, $d,b$, the base matrix compounded to create the configuration $\mathcal{C}_{1}$ will be larger than the one used in $\mathcal{C}_{1}\oplus\mathcal{C}_{2}\oplus\mathcal{C}_{3}$, $\dim(A)_{\mathcal{C}_{1}}>\dim(A)_{\mathcal{C}_{1}\oplus\mathcal{C}_{2}\oplus\mathcal{C}_{3}}$ due to the dimension matching requirements. Therefore as the number of terms in the direct sum decreases along with the compound order, the number of trainable parameters is assumed to increase. One could of course restrict the definition $\mathcal{P}(\mathcal{C})^{\oplus}$ with a fixed base matrix size for all elements, and hence fixed number of parameters, but this may provide a bias in a different direction. As such, we keep the definition flexible and the implication of dimensions will be clear from context through the text.
Next, we have $O\in\{\texttt{comp},\texttt{max},\texttt{avg}\}$, defined as one of the dimensionality-expanding operations on minors from  (13). comp refers to the usual determinant operation on minors. Orthogonality in the adapter matrices is regulated by the binary configuration parameter $\gamma\in\{0,1\}$, with $\gamma=0$ if orthogonality is enforced and $\gamma=1$ otherwise. $\gamma=0$ ensures the transformation preserves the norm and angles of the input feature vectors within the model.
@@
4.4 Quantum native fine-tuning
In this section we briefly describe the quantum-native implementation of our adapters directly on a quantum computer. We expand on this discussion and detail relevant terminology in Appendix A.1. We note that in parallel to our work [24] proposed some ideas for incorporating quantum circuits in fine-tuning pipelines. However, the authors put forward the hardware-efficient family of circuits for fine-tuning, which are known to be problematic due to their problem-agnostic behavior. Their quantum applications have been shown to exhibit barren plateaus [25]. Furthermore, these circuits are not efficiently classically simulatable as the size of the pre-trained matrix, $d$, grows, so no general quantum-inspired approach is possible. In contrast, our compound adapters are both efficiently classically simulatable when the maximum compound rank, $K$, is kept small, and also may provide quantum advantage to implement on quantum computers when both $d$ and $K$, become large. Nevertheless, both the work of [24] and ours demonstrates the feasibility of using quantum and quantum-inspired machine learning for fine-tuning large learning models.
Now, compound adapters can be implemented on quantum computers using circuits composed of quantum gates known or Reconfigurable Beam Splitter (RBS) or their generalization into Fermionic Beam Splitter (FBS) gates. It can be shown, that if complete circuits composed of such gates, act on data encoded within the Hamming-weights of computational basis states the effective action on the input data can be exactly represented by compound matrices. For example, if we have a data vector $\boldsymbol{x}\in\mathbb{R}^{\binom{n}{2}}$ encoded in the amplitudes of the quantum state, $|\psi(\boldsymbol{x})\rangle=\frac{1}{||\boldsymbol{x}||}\sum_{e_{k}\in\text{Hamming-weight}^{n}_{2}}x_{e_{k}}|e_{k}\rangle$ where $e_{k}$ is a bitstring over $n$ (qu)bits with exactly $2$ ones (and $n-2$ zeros, e.g. $0101,1010,1001,0011,1100,0110$ for $n=4$), and we act with certain connectivities of FBS gates on this state, the effective action on the vector is exactly that of the second-order compound, $\mathcal{C}_{2}=A^{(2)}$ [10].
Combined with a suitable tomographic procedure, one may design circuits which can implement all of the above features we found useful - orthogonality, block-sharing and different compound orders. Furthermore, we could also design and use quantum circuits for compound adapters with much larger $K$ than is possible to easily implement in a quantum-inspired way. This ability could open the door to even further compressed fine-tuning in the future.
5 Results and Analysis
Our experiments demonstrate the effectiveness of compound adapters in achieving significant parameter efficiency while maintaining competitive performance across various GLUE benchmark tasks. In this section, we present an analysis of the trade-offs between parameter count and model accuracy, the combined impact of orthogonality and component-wise performance differences.
5.1 Compound Adapters on GLUE dataset
We begin by comparing the best performing compound configuration on the GLUE dataset with the state of the art peft methods in Table 1. This corresponds to the configuration $(\mathcal{C}_{1}\oplus\mathcal{C}_{2},\texttt{comp},b=3,\gamma=0,\beta=0)$, in other words - enforcing orthogonality without block-share over $b=3$ blocks. We see from the Table that, compared to LoRA, we have $\times4$ fewer parameters with an average (relative) performance drop of only $2.9\%$, with 26$\approx$ 26$\times$ 26 × and 25$\approx$ 25$\times$ 25 × fewer parameters than OFT and BOFT with only a $4.3\%$ and $4.4\%$ relative performance drop respectively. These figures are computed relative to the average performance over all GLUE datasets.
Table 2: Summary of configurations with their respective parameter counts and accuracies on the STS-B dataset. We use 3 blocks across all configurations. The best configurations are in bold. If the base matrix is $A$ then $A^{(1)}=:\mathcal{C}_{1}$, $A^{(2)}=:\mathcal{C}_{2}$ and $A^{(3)}=:\mathcal{C}_{3}$. Thus, choosing a configuration with higher Hamming-weight leads to reducing the number of parameters as the size of base matrix $A$ shrinks.
Configuration Base matrix Params Accuracy (%)
$\mathcal{C}_{1}\equiv\text{OFT}$ $\quad A\in\mathbb{R}^{256\times 256}$ $1,\!770,\!241$ $\mathbf{91.68}$
$\mathcal{C}_{2}$ $A\in\mathbb{R}^{23\times 23}$ $38,\!401$ $40.57$
$\mathcal{C}_{3}$ $A\in\mathbb{R}^{12\times 12}$ $16,\!321$ $42.20$
$\mathcal{C}_{1}\oplus\mathcal{C}_{2}$ $A\in\mathbb{R}^{22\times 22}$ $33,\!217$ $\mathbf{88.85}$
$\mathcal{C}_{1}\oplus\mathcal{C}_{3}$ $A\in\mathbb{R}^{12\times 12}$ $16,\!321$ $\mathbf{88.53}$
$\mathcal{C}_{2}\oplus\mathcal{C}_{3}$ $A\in\mathbb{R}^{11\times 11}$ $13,\!057$ $40.60$
$\mathcal{C}_{1}\oplus\mathcal{C}_{2}\oplus\mathcal{C}_{3}$ $A\in\mathbb{R}^{11\times 11}$ $13,\!057$ $\mathbf{88.48}$
Increasing parameters:
From Table 2 we can see two features of our adapters. First, the hyper compression offered by the combinatorial compounding operation, does not allow a large flexibility in changing the number of trainable parameters. Once a non-trivial compound matrix has been added to the adapter (i.e. of greater order than compound $1$ which, when orthogonal, is equivalent to OFT), the parameter count reduces dramatically. This is due to the fixed size of the pre-trained layer, and the large relative size of high-order compounds to lower ones. We provide a visualization of this in the Appendix A.2.1. To address this, we can increase the parameter count monotonically by multiplying several compound adapters. This is a general concept applicable to both additive or multiplicative adapters. For example, in Table 3 we demonstrate that using $4$ multplicative adapters can improve the performance across all GLUE datasets. We only show the best four for compactness.
Table 3: Increasing parameter count in compound adapters. Trainable parameter count can be naturally increased by multiplying successive adapters, leading to performance boosts. Here we compare a single adapter, $\Delta W_{Q}$ versus four, $\prod_{k=1}^{4}(\Delta W^{k}_{Q})$. We show the four GLUE datasets with the largest improvement.
Method # Params CoLA RTE MRPC STS-B
Full Fine-tuning 184M 69.19 83.75 89.46 91.60
$\text{LoRA}_{r=8}$ 1.33M 69.82 85.20 89.95 91.60
$\text{OFT}_{b=16}$ 0.79M 73.91 87.36 92.16 91.91
$\text{BOFT}_{m=2,b=8}$ 0.75M 72.95 88.81 92.40 91.92
$\text{Compound}_{\mathcal{C}_{1}\oplus\mathcal{C}_{2}}$ 0.03M 64.57 81.22 87.99 90.16
$\text{Compound}_{4\times(\mathcal{C}_{1}\oplus\mathcal{C}_{2})}$ 0.14M 65.83 80.50 86.27 91.44
The second observation from Table 2 is the first part of our ablation study. It is clear from these results that our compound adapters cannot be successful without the inclusion of the first order compound - the base matrix, $A=:\mathcal{C}_{1}$, itself. We hypothesise this is due to the difficulty of gradient flow through the determinant operation to the parameters in $A$, when $A$ itself is not included. Also note, it is only the configuration, $\mathcal{C}_{1}$ which is equivalent to OFT, as in this case a single compound matrix (i.e. the trivial one) fills the entire block.
Impact of orthogonality:
The second ablation study we conduct is the impact of orthogonality on the compound adapters. Like the inclusion of the first order compound, we also find including orthogonality is critical for the success of the compound adapters. The possible reason for this is the preservation of orthogonality by determinants, which is reinforced when we study the removal of true compounding - i.e. computing determinants of minors - replaced with other combinatorial operations, such as max and avg (we conduct this ablation study in Appendix A.2.3). Even poorly performing compound configurations, such as those lacking the first order compound, see a significant performance boost when orthogonality is enforced. Finally, we note we show only the impact of orthogonality for multiplicative adapters. One could also consider compound adapters in an additive form (similar to LoRA), which we leave to future work.
Figure 3: Impact of enforcing orthogonality in adapters, for different compound configurations.
5.2 Compound Adapters on VTAB-1K dataset
We then compare the best performing compound configuration on the VTAB subset that we chose. We also reduce the number of examples in each dataset to create VTAB1k [14] where $1000$ random labelled datapoints are used for training an validation, but the final accuracies we show are computed on the entire original VTAB test dataset. Similar to the configuration used on GLUE benchmark, we use $(\mathcal{C}_{1}\oplus\mathcal{C}_{2},\texttt{comp},b=3,\gamma=0,\beta=0)$.
We observe that, compared to $\text{LoRA}_{r=4}$, our method has 13.6$\approx$ 13.6$\times$ 13.6 × fewer parameters while achieving an average relative performance drop of only 0.2%. Similarly, our approach requires 16.2$\approx$ 16.2$\times$ 16.2 × and 15.3$\approx$ 15.3$\times$ 15.3 × fewer parameters than $\text{OFT}_{b=16}$ and $\text{BOFT}_{m=2,b=8}$, with a 0.5% and 0.8% relative performance drop, respectively. These figures are computed relative to the average accuracy across all VTAB tasks in our subset. Interestingly, in contrast with the other datasets across vision and NLP we test, CIFRAR100 stands out as having significantly increased accuracy relative to other methods, on the order of $10\%$.
Table 4: Results on a subset of the VTAB1k benchmark, fine-tuning the pre-trained DINOv2-large model. # Params denotes the number of trainable parameters.Our method is evaluated with the best configuration, $\mathcal{C^{\prime}}=\mathcal{C}_{1}\oplus\mathcal{C}_{2}$, where orthogonality is enforced ($\gamma=0$), parameter sharing across blocks is disabled ($\beta=0$), and the number of blocks is set to $b=3$.
Method # Params (M) CIFAR100 Pets SVHN Resisc45 DMLab
Full Finetuning 304.4M 67.6 93.7 92.8 90.9 58.1
$\text{LoRA}_{r=4}$ 1.77M 77.2 94.8 94.7 91.4 58.1
$\text{OFT}_{b=16}$ 2.10M 77.7 94.7 92.9 91.5 60.5
$\text{BOFT}_{m=2,b=8}$ 1.99M 78.1 95.0 93.0 91.6 61.4
$\text{Compound}_{\mathcal{C}_{1}\oplus\mathcal{C}_{2}}$ 0.13M 87.5 94.04 89.98 88.79 54.74
6 Conclusion
This work presents a new paradigm for parameter-efficient fine-tuning, leveraging quantum-inspired principles to construct effective adapters with minimal additional parameters. Our results indicate that compound matrix-based adapters can serve as a promising alternative to existing PEFT methods (encompassing them in some cases), achieving substantial parameter reduction while maintaining strong performance across a range of NLP and vision tasks.
Our experiments reveal that maintaining orthogonality and selecting an appropriate compound matrix configuration are crucial for preserving model performance. Furthermore, our method’s adaptability across different benchmarks underscores its generalizability and potential for broader applications in resource-constrained environments.
Future work will explore extending these ideas to more complex architectures, further optimizing adapter design, and investigating potential quantum speedups for compound matrix operations. By bridging quantum-inspired techniques with deep learning, we hope to advance the field of efficient fine-tuning and enable scalable adaptation of large foundation models in practical settings.
Appendix A Appendix
A.1 Quantum Implementation
Our adapters, can be implemented efficiently on quantum hardware using fixed Hamming-weight encoders and Hamming-weight preserving circuits. In this section, we detail their implementation on quantum hardware.
A.1.1 RBS gate
@@
$\mathsf{RBS}_{ij}(\theta)=\begin{bmatrix}1&0&0&0\\ 0&\cos(\theta)&\sin(\theta)&0\\ 0&-\sin(\theta)&\cos(\theta)&0\\ 0&0&0&1\end{bmatrix}$
This is a Hamming-weight-preserving gate which is easy to implement on many quantum devices with compilations needing upto $2$ CNOT gates with a Pauli basis native gate set. Another Hamming-weight-preserving gate known as Fermionic Beam Splitter ($FBS$) gate which is a generalisation of $RBS$ gate could also be used to implement Hamming-weight-preserving circuits. The application of a FBS between the qubits $i$ and $j$ , $\mathsf{FBS}_{ij}(\theta)$ , acts as $RBS_{ij}(\theta)$ if the parity of the qubits between $i$ and $j$ is even, and is the conjugate gate $RBS_{i,j}(-\theta)$ otherwise. Therefore, in the case of unary inputs or nearest neighbour connectivity, $FBS$ and $RBS$ gates behave identically. The $FBS_{ij}$ is a non local gate that can be implemented using an RBS gate together with $\mathcal{O}(|i-j|)$ additional two qubit parity gates with a circuit of depth $\mathcal{O}(\log(|i-j|))$. We leave the discussion of quantum adapters using other Hamming-weight preserving (or more specifically particle number preserving) schemes based on Linear Optics [26] circuits for future work.
A.1.2 Loaders
@@
Unary encoding [27, 11] is an amplitude encoding scheme that loads data into the amplitudes of computational basis states where each basis state has a Hamming-weight of 1. It uses $d$ qubits to encode a $d$-dimensional vector. Efficient quantum data encoders using $\mathcal{O}(d)$ two-qubit gates and $\mathcal{O}(\log d)$ depth are known in the unary basis as shown in Fig 4.
Figure 4: A Unary loader. Dots and dashes denote parameterised RBS gates. Figure from  [12]
Fixed Hamming-weight (Hamming-weight-k)  [28] encoding is an amplitude encoding scheme that loads a data vector into a subspace of fixed Hamming-weight $k$. It uses $n$ qubits to encode a data vector of size $d=$(${n}\choose{k}$ , with $n\in\mathcal{O}(kd^{1/k})$. The circuit is constructed using a sequence of controlled (RBS) gates. The total CNOT-gate count for Hamming-weight-k encoding is $\mathcal{O}(kd)$ [28]. This type of encoding is an intermediate regime between unary and binary encodings.
@@
$\text{Total CNOT count}=\sum_{k=1}^{K}\mathcal{O}\left(k\binom{n}{k}\right))\leq\mathcal{O}(d\log d),\ \text{where}\ d={{n}\choose{K}}$ (15)
A.1.3 Layers
Let $G(i,j,\theta)$ denote the Givens rotation applied to the $i$-th and $j$-th unary basis vector, i.e. $e_{i}$ and $e_{j}$, $\boldsymbol{\theta}$ a vector of angles, and $\mathcal{T}$ is a list of triplets $(i,j,m)$. The Hamming-weight-preserving layer is defined by:
$U(\boldsymbol{\theta})=\prod_{(i,j,m)\in\mathcal{T}}\text{RBS}_{ij}(\theta_{m}).$
It acts as $U(\boldsymbol{\theta})|\boldsymbol{x}\rangle=W|\boldsymbol{x}\rangle$ where $W=\prod_{(i,j,m)\in\mathcal{T}}G(i,j,\theta_{m})$.
@@ A.1.3 Layers
Figure 5: Hamming-weight preserving layers. Dots and dashes denote parameterised RBS gates. Figure from  [12].
There are different circuits for $U(\boldsymbol{\theta})$, highlighted in Figure 5. The Pyramid architecture, as described in  [11], consists of $n(n-1)/2$ RBS gates arranged in a pyramid-like structure and has a linear depth. This architecture allows for the representation of all possible orthogonal matrices of size $n\times n$. The Butterfly architecture, which was proposed in [29], in uses logarithmic depth circuits with a linear number of gates to implement a quantum orthogonal layer. This architecture, classical Cooley–Tukey algorithm used for Fast Fourier Transform, requires all-to-all connectivity in the hardware layout.
A.1.4 Quantum Implementation
We can use these tools to construct quantum native implementation of our adapters as shown in figure 6. The block diagonal structure of our adapters imply that the adapters can be implemented via separate quantum circuits. For example in figure 6(a), a 4 block $\mathcal{C}_{1}$ adapter can be implemented via 4 quantum circuits, each with Hamming-weight-1 loaders, a Hamming-weight-preserving layer and suitable measurements. Enforcing block share in this setting would imply the circuit layers sharing the same parameter values, however, the loaders still ought to be different. Similarly in figure 6(b), we use 2 quantum circuits each with Hamming-weight-1, Hamming-weight-2 and Hamming-weight-3 loaders stacked one after another. Note that as specified in the binary encoders of  [28], we would need parameterised controlled $R_{Y}$ gates between each loader to enable sequential stacking.
Figure 6: Quantum Implementation of Adapters. Each compound adapter has an efficient quantum implementation using fixed Hamming-weight encoders and Hamming-weight preserving layers. Trailing dimensions are padded with an identity matrix. The figure shows quantum circuits for a) $\mathcal{C}_{1}$, $b=4$ blocks, which uses only Hamming-weight 1 loaders and b) $\mathcal{C}_{1}\oplus\mathcal{C}_{2}\oplus\mathcal{C}_{3}$, $b=2$ blocks which uses upto Hamming-weight 3 loaders.
A.2 Ablation studies on STS-B dataset
To further understand the impact of different configuration setups, we run ablation studies on one dataset STS-B.
A.2.1 Compound Configurations
As illustrated in Figure 7, we explore how different configurations of compound adapters perform on the STS-B dataset - an illustration of Table 2 in the main text. We note that the presence of $\mathcal{C}_{1}$ adapter with higher orders show the best performance while giving significant parameter reductions compared to only having higher order adapters ($\mathcal{C}_{2}$ or $\mathcal{C}_{3}$).

### v2

@@ Appendix C Adapter Configurations (Extended)
Here we elaborate on the different possible configurations of a QuIC Adapter. Our experimentation focused on different combinations of compound matrices based on Hamming-weights, the types of operations applied to these compounds, the enforcement of orthogonality, and the strategy for parameter sharing across adapter blocks.
@@
$\mathcal{C}:=\{\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{C}_{3}\},\\ \mathcal{P}(\mathcal{C})^{\oplus 3}:=\{\mathcal{C}_{1},\mathcal{C}_{2},\mathcal{C}_{3},\mathcal{C}_{1}\oplus\mathcal{C}_{2},\mathcal{C}_{2}\oplus\mathcal{C}_{3},\mathcal{C}_{2}\oplus\mathcal{C}_{3},\mathcal{C}_{1}\oplus\mathcal{C}_{2}\oplus\mathcal{C}_{3}\}$ (10)
Note that this notation is slightly obfuscating. Given a fixed pre-trained matrix and block size, $d,b$, and two different configurations both containing the base matrix, e.g. $\mathcal{C}_{1}$ and $\mathcal{C}_{1}\oplus\mathcal{C}_{2}\oplus\mathcal{C}_{3}$. The base matrix compounded to construct the former configuration will be larger (and hence have more trainable parameters) than the one used to create the latter, in other words $\dim(A)_{\mathcal{C}_{1}}>\dim(A)_{\mathcal{C}_{1}\oplus\mathcal{C}_{2}\oplus\mathcal{C}_{3}}$ due to the dimension matching requirements. Therefore as the number of terms in the direct sum decreases along with the compound order, the number of trainable parameters is assumed to increase. One could of course restrict the definition $\mathcal{P}(\mathcal{C})^{\oplus}$ with a fixed base matrix size for all elements, and hence fixed number of parameters, but this may provide a bias in a different direction. As such, we keep the definition flexible and the implication of dimensions will be clear from context through the text.
Next, we have $O\in\{\texttt{comp},\texttt{max},\texttt{avg}\}$, defined as one of the dimensionality-expanding operations on minors from above, or ‘compounding’ - comp - which refers to the usual determinant operation on minors. Orthogonality in the adapter matrices is regulated by the binary configuration parameter $\gamma\in\{0,1\}$, with $\gamma=0$ if orthogonality is enforced and $\gamma=1$ otherwise. $\gamma=0$ ensures the transformation preserves the norm and angles of the input feature vectors within the model.
@@
Appendix D Quantum Implementation
[…]
D.1 Reconfigurable Beam Splitter gates
@@
$\text{RBS}_{ij}(\theta)=\begin{bmatrix}1&0&0&0\\ 0&\cos(\theta)&\sin(\theta)&0\\ 0&-\sin(\theta)&\cos(\theta)&0\\ 0&0&0&1\end{bmatrix}$
This is a Hamming-weight-preserving gate which is easy to implement on many quantum devices with compilations needing upto $2$ CNOT gates with a pauli basis native gate set. Another Hamming-weight-preserving gate known as Fermionic Beam Splitter ($FBS$) gate which is a generalisation of $RBS$ gate could also be used to implement Hamming-weight-preserving circuits. The application of a $FBS$ between the qubits $i$ and $j$ , $FBS_{ij}(\theta)$ , acts as $RBS_{ij}(\theta)$ if the parity of the qubits between $i$ and $j$ is even, and is the conjugate gate $RBS_{i,j}(-\theta)$ otherwise. Therefore, in the case of unary inputs or nearest neighbour connectivity, $FBS$ and $RBS$ gates behave identically. The $FBS_{ij}$ is a non local gate that can be implemented using an RBS gate together with $\mathcal{O}(|i-j|)$ additional two qubit parity gates with a circuit of depth $\mathcal{O}(\log(|i-j|))$. We leave the discussion of quantum adapters using other Hamming-weight-preserving modalities like Linear Optics circuits for future work.
D.2 Loaders
@@
[…]
Unary encoding [Johri et al., 2021, Landman et al., 2022] is an amplitude encoding scheme that loads data into the amplitudes of computational basis states where each basis state has a Hamming-weight of 1. It uses $d$ qubits to encode a $d$-dimensional vector. Efficient quantum data encoders using $\mathcal{O}(d)$ two-qubit gates and $\mathcal{O}(\log d)$ depth are known in the unary basis as shown in Fig 5.
[…]
Figure 6: A fixed Hamming-weight encoder. Figure shows loading Hamming-weight-22 subspace ($k=2$) in 66 qubits. Blue and green denote input and output respectively, violet denotes controlled operation. Figure from  [Farias et al., 2024].
@@
$\text{Total CNOT count}=\sum_{k=1}^{K}\mathcal{O}\left(k\binom{n}{k}\right)\leq\mathcal{O}(d\log d),\ \text{where}\ d={{n}\choose{K}}$ (11)
D.3 Layers
Let $G(i,j,\theta)$ denote the Givens rotation applied to the $i$-th and $j$-th unary basis vector, i.e. $e_{i}$ and $e_{j}$, $\mathbf{\theta}$ a vector of angles, and $\mathcal{T}$ is a list of triplets $(i,j,m)$. The Hamming-weight-preserving layer is defined by:
$U(\mathbf{\theta})=\prod_{(i,j,m)\in\mathcal{T}}\text{RBS}_{ij}(\theta_{m}).$
[…]
@@ D.3 Layers
Figure 7: Hamming-weight preserving layers. Dots and dashes denote parameterised RBS gates. Figure from  [Cherrat et al., 2023].
There are different circuits for $U(\mathbf{\theta})$, highlighted in Figure 7. The Pyramid architecture, as described in  [Landman et al., 2022], consists of n(n1)/2n(n-1)/2 RBS gates arranged in a pyramid-like structure and has a linear depth. This architecture allows for the representation of all possible orthogonal matrices of size $n\times n$. The Butterfly architecture, which was proposed in [Cherrat et al., 2024], in uses logarithmic depth circuits with a linear number of gates to implement a quantum orthogonal layer. This architecture, classical Cooley–Tukey algorithm used for Fast Fourier Transform, requires all-to-all connectivity in the hardware layout.
D.3.1 Quantum Implementation
We can use these tools to construct quantum native implementation of our adapters as shown in figure 8. The block diagonal structure of our adapters imply that the adapters can be implemented via separate quantum circuits. For example in figure 8(a), a 4 block $\mathcal{C}_{1}$ adapter can be implemented via 4 quantum circuits, each with Hamming-weight-1 loaders, a Hamming-weight-preserving layer and suitable measurements. Enforcing block share in this setting would imply the circuit layers sharing the same parameter values, however, the loaders still ought to be different. Similarly in figure 8(b), we use 2 quantum circuits each with Hamming-weight-1, Hamming-weight-2 and Hamming-weight-3 loaders stacked one after another. Note that as specified in the binary encoders of  [Farias et al., 2024], we would need parameterised $R_{Y}$ gates between each loader to enable sequential stacking.
Figure 8: Quantum Implementation of Adapters. Each QuIC Adapter has an efficient quantum implementation using fixed Hamming-weight encoders and Hamming-weight preserving layers. Trailing dimensions are padded with an identity matrix. The figure shows quantum circuits for a) $\mathcal{C}_{1}$, $b=4$ blocks, which uses only Hamming-weight 1 loaders and b) $\mathcal{C}_{1}\oplus\mathcal{C}_{2}\oplus\mathcal{C}_{3}$, $b=2$ blocks which uses upto Hamming-weight 3 loaders.
D.4 Ablation studies on STS-B dataset